OpenStudy (anonymous):

Hi could someone please check this for me :) If x^2 = y^3 then x^(2/3) = y?

OpenStudy (loser66):

yup

OpenStudy (anonymous):

Let's see, You're pretty much dividing the exponents to x^2 = y^3 would be equal to x^2/3 = y